OpenStudy (anonymous):

Given the function g(x) = −4x + 5, find g(−4).

OpenStudy (anonymous):

21

OpenStudy (anonymous):

g(-4) = -4(-4) + 5 g(-4) = 16 +5 g(-4) = 21

OpenStudy (anonymous):

g(-4)=-4 x -4 +5= = +16+5 = 21

OpenStudy (anonymous):

21

OpenStudy (anonymous):

thankyou everybody for your help